OpenStudy (anonymous):

Simplify 2^4 + 3^2 14 24 25 64

OpenStudy (shamim):

2^2=?

OpenStudy (anonymous):

its 25

OpenStudy (anonymous):

the power by 4/2

OpenStudy (shamim):

u know 2^2=2*2=4 right?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

tell me plz 2^3=?

OpenStudy (anonymous):

8?

OpenStudy (shamim):

good

OpenStudy (shamim):

then tell me 2^4=?

OpenStudy (anonymous):

16

OpenStudy (shamim):

u r right

OpenStudy (shamim):

3^2=?

OpenStudy (anonymous):

Oh ok I got it now thanks so much!

OpenStudy (shamim):

u r most welcome